A nut mixture consists of almonds and cashews. Almonds are $4.98 per pound, and cashews are $6.98 per pound. How many pounds of each type of nut should be mixed to produce a 16 pounds selling for $5.73 per pound.
Answer by edjones(8007)

Let x=lbs of almonds
16-x=lbs of cashews
4.98x+6.98(16-x)=5.73*16
4.98x+111.68-6.98x=91.68
-2x+111.68=91.68add 2x and subtract 91.68 from each side: 2x=20
divide 2 into each side: x=10
Check:
(4.98*10)+111.68-69.8=91.68
49.8+111.68-69.8=91.6
91.68=91.68
Ed